### v1

# VisuRiddles: Fine-grained Perception is a Primary Bottleneck for Multimodal Large Language Models in Abstract Visual Reasoning

Abstract
Recent strides in multimodal large language models (MLLMs) have significantly advanced their performance in many reasoning tasks. However, Abstract Visual Reasoning (AVR) remains a critical challenge, primarily due to limitations in perceiving abstract graphics. To tackle this issue, we investigate the bottlenecks in current MLLMs and synthesize training data to improve their abstract visual perception. First, we propose VisuRiddles, a benchmark for AVR, featuring tasks meticulously constructed to assess models’ reasoning capacities across five core dimensions and two high-level reasoning categories. Second, we introduce the Perceptual Riddle Synthesizer (PRS), an automated framework for generating riddles with fine-grained perceptual descriptions. PRS not only generates valuable training data for abstract graphics but also provides fine-grained perceptual description, crucially allowing for supervision over intermediate reasoning stages and thereby improving both training efficacy and model interpretability. Our extensive experimental results on VisuRiddles empirically validate that fine-grained visual perception is the principal bottleneck and our synthesis framework markedly enhances the performance of contemporary MLLMs on these challenging tasks. Our code and dataset will be released at https://github.com/yh-hust/VisuRiddles
footnotetext: Project Lead
Figure 1: Performance and analysis of MLLMs on AVR. (a) Most advanced MLLMs achieve limited accuracy on VisuRiddles, often close to random choice and far below human performance. (b) Model responses to abstract graphics and their perceptual descriptions show that once equipped with perceptual understanding, MLLMs can succeed in AVR task.
1 Introduction
Visual reasoning is a core component of human intelligence [1]. Enhancing this capacity in Multimodal Large Language Models (MLLMs) is key to bridging the gap to human-level performance [10, 11]. Recent advancements [4, 5, 6, 75, 9, 8, 7] in MLLMs have made significant progress in various general visual tasks, but still struggle with Abstract Visual Reasoning (AVR). As shown in Fig. 1(a), even advanced models such as Gemini2.5 Pro continue to face significant challenges in this task, showing a substantial gap compared to human performance. While such AVR tasks are relatively easy for humans, even the most advanced MLLMs struggle to solve them. It seems unexpected that models with strong performance in general understanding and reasoning tasks still fall short in handling such reasoning challenges. In fact, the core issue lies not in reasoning but in the fine-grained perception of complex visual structures. Unlike humans, who naturally process fine-grained visual structural cues, MLLMs often lack this perceptual ability needed for high-level abstract reasoning. As shown in Fig. 1(b), once the abstract graphics is reformulated into fine-grained perceptual descriptions, the model can produce accurate responses. This highlights that enhancing fine-grained perceptual capacity is crucial to unlocking MLLMs’ potential in AVR.
The significance of AVR tasks derives not only from their inherent complexity but also from their role in revealing the fundamental nature of human perception and cognition [62]. Therefore, abstract graphics serves as a prototypical and challenging paradigm for evaluating the reasoning capabilities of MLLMs [2, 3]. Cognitive research suggests that AVR requires the integration of multiple visual dimensions within abstract graphics, including numerosity, attributes, style, position, and spatial relations, which correspond respectively to sensitivity to quantity, object features, stylistic variation, relative placement, and spatial arrangement [63, 64, 65, 66]. Visual abstraction abilities across these five dimensions form the foundation of visual pattern understanding and serve as indispensable cognitive supports for high-level visual logical reasoning. A deficiency in perceiving the visual information of these dimensions can significantly limit a MLLM’s performance on AVR tasks.
Although solving abstract graphics problems involves multiple cognitive dimensions, benchmarks explicitly designed to evaluate them along these dimensions remain limited. To address this gap, we introduce VisuRiddles, a benchmark for evaluating AVR based on the five core dimensions and two high-level reasoning categories. VisuRiddles enables a more comprehensive assessment of abstract visual reasoning. However, most existing datasets offer only question–answer pairs, lacking the perceptual process supervision needed to bridge visual input and reasoning. To tackle this challenge, we further introduce an automated data synthesis framework that generates aligned samples with fine-grained perceptual descriptions, enabling granular supervision across the full perception-to-reasoning pipeline for more systematic learning and improved generalization.
Our contributions can be summarized as follows:
  • We introduce VisuRiddles, a comprehensive benchmark for AVR, enabling multi-dimensional evaluation of MLLMs across diverse reasoning dimensions.
  • We present the Perceptual Riddle Synthesizer, an automated framework designed to address the lack of intermediate supervision by producing structured AVR riddles enriched with fine-grained perceptual descriptions.
  • We conduct extensive experiments to demonstrate our key findings, including the reasoning limitations of current MLLMs, the effectiveness of our synthesized riddles, and the effective approach to addressing abstract visual problems.
2 Related Work
2.1 Multimodal Benchmarks.
High-quality benchmarks are essential for driving progress in MLLMs. Earlier evaluation effort predominantly centered on basic visual comprehension tasks, leading to the development of extraction-oriented benchmarks [12, 13, 14, 15, 76, 76, 77, 16]. These benchmarks primarily assess a model’s ability to identify and align explicit visual elements. However, they depend heavily on surface-level cues, offering little assessment of an MLLM’s capacity for abstraction and logical reasoning. Therefore, researchers have expanded their focus on more holistic evaluation frameworks, proposing general-purpose benchmarks [17, 18, 19, 20, 78]. These benchmarks span multiple tasks, including visual understanding, mathematical reasoning, and knowledge-based QA, and are designed to assess the generalist capabilities of MLLMs. Although general-purpose benchmarks span a wide range of tasks, they tend to emphasize knowledge-based evaluation over visual logical reasoning. Moreover, due to their strong semantic dominance, MLLMs may directly rely on their inherent knowledge, making it difficult to disentangle and evaluate models’ true visual reasoning capabilities. To address these limitations, certain studies have shifted toward evaluating visual logical reasoning, leading to logic-oriented benchmarks such as PGM [24], RAVEN [21], I-RAVEN [22], RAVEN-FAIR [23], CVR [25], Raven IQ [26], ConceptARC [27], RPMs [28], MaRs-VQA [2] MathVista [29], GeoEval [30], MathVision[31], MathVerse [32], CMMath [33], OCR-Reasoning [78] and MV-MATH [34]. These benchmarks use structured diagrams, mathematical forms, and spatial layouts to assess analogical reasoning, compositionality, and pattern induction, signaling a move toward higher-level AVR. However, these benchmarks still exhibit notable limitations, particularly in their partial reliance on external knowledge and lack breadth in reasoning coverage.
2.2 Multimodal Large Language Models for Visual Reasoning.
Advanced commercial LLMs, including the GPT [35], Gemini [36], and Claude [37] series, have exhibited impressive capabilities in multimodal understanding and generation. These advances have stimulated increasing interest and systematic investigation into MLLMs from both academia and industry. Early representative general-purpose MLLMs, such as LLaVA [38], Instruct-BLIP [39], Qwen-VL [40], Intern-VL [41], and mPLUG-Owl [42], perform well on general multimodal tasks but exhibit limitations in fine-grained visual perception. Subsequently, DeepSeek-VL [43], Monkey [8], InternVL 1.5 [44], TextMonkey [45], InternLM-XComposer2-4KHD [46], Qwen2-VL [47], and mPLUG-Owl2 [48] have introduced high-resolution visual perception mechanisms, establishing a more robust foundation for fine-grained visual understanding and reasoning. However, such improvements primarily expand the perceptual scope of MLLMs without substantially enhancing their reasoning capabilities. Therefore, inference-time scaling methods including LLaVA-CoT[49], BBA [50], R-CoT [51], and RedStar [52] have introduced and innovated upon inference-time scaling techniques such as chain-of-thought (CoT), significantly enhancing the visual reasoning capabilities of MLLMs. More recently, following the advances of Kimi K1.5 [53] and DeepSeek-R1 [54], a series of studies, such as R1-OneVision [3], LMM-R1 [55], MM-EUREKA [56], R1-V [57], Visual-RFT [58], VisualPRM [59], OThink-MR1 [60], and VLM-R1 [61], have employed reinforcement learning to develop adaptive reasoning policies and improve generalization in complex visual reasoning tasks.
Figure 2: Representative examples from VisuRiddles. The benchmark includes eight reasoning categories, designed to comprehensively evaluate diverse reasoning capabilities of MLLMs.
3 Method
In this section, we first describe the construction process of VisuRiddles, following a structured pipeline consisting of data collection, cleaning, and consolidation. We then summarize key benchmark statistics—including question and image counts, as well as answer distribution in Sec. 3.1. Subsequently, we introduce the Riddles Synthesis Framework, an automated approach for generating structurally diverse and perceptually annotated samples tailored for AVR in Sec. 3.2.
3.1 VisuRiddles Benchmark
Existing benchmarks provide limited coverage of the core reasoning dimensions required for AVR, To address this, VisuRiddles explicitly evaluates models across five key dimensions: a) Numerosity assesses the model’s ability to perceive and reason about quantity and distribution; b) Attribute evaluates the understanding of intrinsic visual features that determine structural semantics; c) Style tests the capacity to identify and generalize transformation-based visual patterns; d) Position reflects the ability to reason over the relative positions and layout of visual elements; e) Spatiality examines the understanding of three-dimensional structures, shape variations, and spatial transformations in abstract scenes. To extend the evaluation beyond low-level perceptual reasoning and address limitations in task diversity, structural complexity, and the inherent uncertainty introduced by single-choice formats, we additionally incorporate two high-level reasoning tasks that require analogical abstract reasoning and consistency-based logical reasoning, as represented by RAVEN and Sudoku, respectively. Finally, VisuRiddles also includes a subset, namely Other of AVR tasks involving planar shape composition and character-based semantic patterns. VisuRiddles provides a unified benchmark spanning basic to high-level reasoning, enabling comprehensive assessment of AVR, with representative samples shown in Fig. 2.
We construct the VisuRiddles benchmark through a structured pipeline comprising Collection, Cleaning, and Consolidation:
Collection. We manually collect a diverse set of visual reasoning problems from Chinese National Civil Service Examination, which form the five core categories of the benchmark. RAVEN and Sudoku samples are generated following the method1111https://github.com/WellyZhang/RAVEN1 and the protocols2222https://gitee.com/mxx11/sudoku2.
Cleaning. After initial collection, we recruited twelve trained annotators to verify and refine the dataset. This process involved identifying and removing duplicated questions, correcting answer key errors, and filtering out incomplete or noisy samples to ensure data consistency and reliability. Quality control included annotator training, multi-pass review, dispute resolution, and cross-validation on overlapping samples.
Consolidation. All cleaned samples were manually translated into English, covering question and answer. Each item was then categorized into its corresponding reasoning type based on expert explanations provided by the Fenbi educational platform. All data were standardized into a unified JSON format, supporting structured access to image content, metadata, and reasoning annotations.
Following the structured pipeline, we initially collected 1,275 abstract visual reasoning problems, each accompanied by expert-level analysis and answer. After careful filtering and quality control, we construct VisuRiddles with 770ok samples in the basic categories and 200 samples in the high-level categories. Key statistics of VisuRiddles are summarized in Tab. 2. For the basic AVR tasks, each question is formulated as a single-choice problem, with answer option distributions of A (25.75%), B (24.75%), C (24.375%), and D (25.125%). In the high-level categories, RAVEN Reasoning and Sudoku Reasoning each account for 10% of the total data, and require models to generate exact symbolic outputs to be considered correct.
Table 1: The statistics of VisuRiddles, including the number questions, average question length (in tokens), number of associated images, and the corresponding answer format.
Type Questions Number Question length Image Number Answer Formats
Sudoku Reasoning 100 48.3 100 Constraint-based grid
RAVEN Reasoning 100 70.7 200 Pattern match index
Numerical Reasoning 250 26.3 250 Single-choice
Stylistic Reasoning 117 26.3 117 Single-choice
Attribute Reasoning 97 27.0 97 Single-choice
Positional Reasoning 111 25.5 111 Single-choice
Spatial Reasoning 156 27.2 156 Single-choice
Other 69 30.3 69 Single-choice
Total 1000 32.5 1100
3.2 Perceptual Riddle Synthesizer
Recent MLLMs demonstrate strong general reasoning, but their limited performance on AVR is largely attributed to insufficient capacity of fine-grained perception. Most existing datasets lack intermediate perceptual description, providing only question-answer pairs. This prevents explicit modeling of the perception-to-reasoning process, leading to black-box inference, weak inductive capacity, and poor generalization. More critically, annotating such perceptual processes requires effort of expert-level staff members, further compounding this problem. To address this, we introduce Perceptual Riddle Synthesizer (PRS), a Riddles Synthesis Framework to provides abstract graphics with aligned perceptual description.
Figure 3 illustrates the pipeline of our PRS framework for AVR task. The pipeline consists of three main stages:
Rule Selection. The synthesis process begins with rule selection. We first extract representative reasoning patterns from real-world standardized tests, such as civil service visual logic exams. These high-level rules are categorized into positional, stylistic, numerical,attribute,spatial,raven and sudoku types, and further decomposed into finer-grained sub-rules, each representing a distinct visual transformation or logical operation (e.g., “simultaneous rotation and revolution”, or “black-white XOR operation”). This decomposition enables modular and interpretable synthesis, allowing each instance to target a specific reasoning skill.
Element Configuration. Once a sub-rule is selected, we proceed to define the visual components in a structured manner. This includes: (i) choosing a layout template (e.g., $3\times 33 × 3$ matrix, progression row, etc.), (ii) selecting or auto-generating icon elements with specific visual properties (e.g., shape, color, direction), and (iii) assigning a background grid. Elements are configured in a rule-consistent way: for example, a “positional” sub-rule may require translation across rows, while a “stylistic” sub-rule may apply logical operations to shape fill. Each configuration ensures the pattern remains visually interpretable and solvable.
Rendering and Annotating.
Each question instance is constructed by applying parameterized transformations under defined constraints: a single correct answer, consistent application of one rule, and plausible distractors. The resulting sequence of visual instances is rendered into final image grids. In parallel, the framework generates structured perceptual description in JSON format, including layout metadata, element-level properties (e.g., object shape, position, state), and reasoning traces. These annotations bridge perception and reasoning, enabling models to explicitly learn visual abstraction rather than relying on black-box pattern matching.
PRS produces a large-scale dataset spanning five core perceptual reasoning types and two high-level structured reasoning tasks. It supports 34 sub-rules across these categories and generates samples using 8 layout and 10k unique graphical icons. Each instance is created through rule-guided visual composition and rendered alongside structured perceptual annotations. The resulting dataset exhibits high diversity in both visual patterns and reasoning logic, as summarized in Tab. 2. It is worth noting that the synthetic tasks are designed to strengthen perceptual abstraction rather than replicate real-world reasoning difficulty. Consequently, their complexity is deliberately kept tractable to focus on providing structured perceptual supervision.
Figure 3: Overview of the PRS framework. (a) A unified pipeline for generating abstract graphics with fine-grained perceptual descriptions across different rule categories. (b) Visualization of synthesized riddles based on positional rule and stylistic rule.
Table 2: Configuration details of PRS across seven reasoning categories.
Rule Sub-rule Layout Template Icon Sample
Numerical 5 2 1 5k 10K
Stylistic 6 3 5 5k 10K
Attribute 2 2 1 5k 10K
Positional 8 4 7 5k 10K
Spatial 5 2 13 5k 10K
Sudoku 1 1 4 9 10K
RAVEN 7 5 5 10k 20K
4 Experiment
In this section, we present extensive experiments on advanced MLLMs on the challenging VisuRiddles benchmark. We first describe the experimental setup for both training and evaluation in Sec. 4.1. Subsequently, we demonstrate a comprehensive evaluation of MLLMs in Sec. 4.2. We further include visualization examples that illustrate model behaviors and failure modes in Sec. 4.3. Finally, we conduct ablation studies to validate the efficacy of our synthesis engine in Sec. 4.4.
4.1 Experimental Setup
Training details. We fine-tune the open-source model Qwen2.5-VL-7B using 80k perceptually annotated samples generated by our Riddles Synthesis Framework. Training is performed with a learning rate of 5e-6, using the Adam optimizer, a batch size of 16, and 8 NVIDIA A800 GPUs.
Evaluation Settings. We evaluate seventeen MLLMs, including eleven open-source and six advanced proprietary models. The open-source group consists of MiniCPM-V 2.6 [7], DeepSeek-VL2 [67], Qwen2.5-VL [68], InternVL2.5 [5] and InternVL2.5-MPO (Mixed Preference Optimization) [69], which is designed to improve reasoning capabilities via reinforcement learning-based optimization. The proprietary group includes GPT-4o (20241120) [70], o1 [71], Claude-3-7-Sonnet (20250219) [72], Claude-3-7-Sonnet-Thinking (20250219) [72], Gemini-2.0-Flash-Thinking (20250121) [73], and Gemini-2.5-Pro (20250325) [74]. Notably, GPT-o1, Claude-3-7-Sonnet-Thinking, and Gemini-2.5-Pro are equipped with specialized “thinking” modes aimed at enhancing multi-step reasoning. To further explore reasoning capabilities, CoT prompting is applied to GPT-4o, Qwen2.5-VL-72B, and InternVL2.5-78B-MPO, selected for their reasoning capacity, model scale, or lack of built-in “thinking” mode. Inference for all models is performed using the vLLM framework with unified decoding configurations. Human reference answers are obtained from the Fenbi Educational Platform, a widely used source for civil service exam preparation.
Table 3: Evaluation of various models on VisuRiddles. The superscript number indicates the number of answer options (e.g., “4” denotes a 4-choice question where one option is correct; $\infty$” represents unlimited possible options). The best results are marked bold and the second results are underlined.
[…]
Human - 61.3 60.9 67.5 67.9 58.8 - - 61.9 63.0
Open-Source MLLMs
Minicpm-V-2.6 8B 22.2 23.1 38.9 23.1 27.4 1.0 11.0 28.2 21.9
InternVL2.5-8B 8B 23.0 29.9 27.8 31.7 28.0 0.0 8.0 31.0 22.7
InternVL2.5-8B-MPO 8B 25.7 29.9 20.0 27.9 27.4 0.0 10.0 31.0 22.4
Deepseekvl2 27B 28.4 29.1 22.2 22.1 24.2 0.0 8.0 18.3 21.0
InternVL2.5-38B 38B 22.6 26.5 28.9 27.9 26.8 1.0 9.0 31.0 21.9
InternVL2.5-38B-MPO 38B 24.5 23.1 33.3 29.8 27.4 0.0 8.0 25.4 22.1
Qwen2.5VL-32B 32B 25.7 23.9 24.4 29.8 28.7 1.0 9.0 25.4 22.1
InternVL2.5-78B 78B 21.8 31.6 30.0 27.9 26.8 0.0 15.0 22.5 22.3
InternVL2.5-78B-MPO 78B 26.8 31.6 25.6 26.9 24.8 0.0 13.0 23.9 22.7
InternVL2.5-78B-MPO(cot) 78B 26.1 24.8 27.8 33.7 21.7 0.0 16.0 21.1 22.2
Qwen2.5VL-72B 72B 24.5 23.1 23.3 29.8 22.3 2.0 12.0 29.6 21.3
Qwen2.5VL-72B(cot) 72B 26.4 25.6 24.4 32.7 26.8 0.0 18.0 31.0 23.7
Commercial products
GPT-4o - 24.5 22.2 28.9 25.0 22.3 0.0 17.0 25.4 21.2
GPT-4o(cot) - 28.7 28.2 34.4 27.9 22.9 0.0 11.0 33.8 23.9
GPT-o1 - 31.0 28.2 33.3 25.0 26.1 0.0 15.0 26.8 24.5
claude-3-7-sonnet - 25.3 23.1 30.0 26.0 25.5 1.0 12.0 28.2 22.0
claude-3-7-sonnet-thinking - 21.1 26.5 41.1 30.8 26.8 9.0 13.0 26.8 23.8
Gemini-2.0-flash-thinking - 17.6 12.8 35.6 26.9 17.2 10.0 11.0 25.4 18.7
Gemini2.5-pro - 26.4 32.5 44.4 29.8 23.6 15.0 14.0 33.8 26.8
Ours
Baseline(Qwen2.5VL-7B) 7B 23.4 28.2 28.9 23.1 21.7 0.0 14.0 23.9 20.9
PRS-Qwen2.5VL-7B 7B 27.2 35.7 41.4 38.3 30.3 60.0 95.0 21.3 43.7
4.2 Main Results
Table 3 presents performance of various MLLMs on VisuRiddles benchmark. We summarize the key observations as follows:
Performance of Open-Source MLLMs. Most open-source MLLMs demonstrate limited performance on VisuRiddles, with accuracies across core categories often close to random choice and consistently failing on the two high-level reasoning tasks, which reveals a substantial gap between current model capabilities and human performance in AVR.
Performance of Commercial Models with Thinking Mode. Several commercial MLLMs with ‘thinking’ mode, such as GPT-o1, Claude-3-7-Sonnet-Thinking, and Gemini2.5-Pro, achieve comparatively improved performance on VisuRiddles. Their advantage is particularly notable in structured tasks like Sudoku and Raven. These results indicate that AVR can benefit from advanced reasoning capabilities. However, the overall performance remains unsatisfactory, suggesting that reasoning enhancements alone are insufficient to address the core challenges of this task.
Effectiveness of Model Scaling and CoT Prompting. Strategies such as increasing model size and applying CoT prompting, though commonly used to enhance reasoning, are ineffective for AVR. Models with larger parameter scales do not consistently outperform smaller ones, and CoT yields only limited improvements on select tasks. These findings indicate that scaling in model parameter or inference time is insufficient to address the core challenges of AVR.
Effectiveness of Our Perception-Enhanced Model. Our perception-enhanced model significantly outperforms other MLLMs across all categories of VisuRiddles, particularly in high-level reasoning tasks, Sudoku reasoning and Raven reasoning, demonstrating the model’s ability to handle structurally complex visual patterns. Importantly, these gains are achieved without changing model architecture or size, but solely through training on data with perceptual description, highlighting that the improvement arises from enhanced perceptual abstraction rather than stronger reasoning capacity. These results affirm the critical role of fine-grained perceptual supervision in abstract visual reasoning and emphasize the need to model intermediate perceptual processes for better performance on structurally complex tasks.
Figure 4: Case study comparing different reasoning strategies on a VisuRiddles example. (a) QVQ-72B reflects a flawed loop under the ”thinking” mode. (b) Qwen2.5-VL-72B with CoT prompting exhibits incorrect perceptual understanding. (c) Our perception-enhanced model generates structured reasoning grounded in accurate perceptual abstraction, yielding the correct answer.
4.3 Visualization Results
To further analyze the reason that MLLMs underperform on AVR task, we examine two widely adopted reasoning enhancement strategies: inference-time scaling and CoT prompting. We conduct experiments on QVQ-72B-Preview (equipped with a ”think” mode), Qwen2.5-VL-72B (with CoT prompting), and our perception-enhanced model Qwen2.5-8B.
Fig. 4 visualizes a representative case of a broadly observed phenomenon uncovered during our experiments. The results reveal clear differences in the reasoning behaviors across these models. The inference-time scaling approach, which leverages ”think”-style prompting without any perceptual grounding, frequently leads to verbose yet logically inconsistent reasoning, often ending in incorrect answers. The CoT prompting, while occasionally yielding the correct answer, fails to guide the model along a valid reasoning path due to inaccurate visual perception. In such cases, correct predictions can be largely attributed to the single-choice format rather than abstract visual understanding. This observation aligns with CoT-based models perform close to random guessing on basic inference tasks, but struggle significantly on high-level inference tasks, shown in Tab. 3. In contrast, our approach leverages fine-grained perception to produce concise, visually grounded reasoning and accurate answers.
These results underscore the importance of perception in abstract visual reasoning. Inference-time scaling and CoT prompting alone cannot compensate for the lack of perceptual grounding, which often leads to incorrect or inconsistent reasoning. In contrast, introducing fine-grained perceptual supervision enables the model to build accurate and visually aligned reasoning paths, supporting reliable performance on complex visual tasks.
4.4 Ablation Study
Ablation Study on Bottlenecks in Abstract Visual Reasoning. To investigate whether the performance gap in abstract visual reasoning stems from perceptual limitations rather than reasoning ability, we conduct an ablation study using our perception-annotated synthetic dataset. We evaluate two representative MLLMs, GPT-4o (closed-source) and Qwen2.5VL-72B (open-source), on a subset of our synthetic dataset under two input settings: (i) raw abstract graphics (V), and (ii) structured perceptual descriptions (P).
The experimental results in Tab. 4 show that both models perform poorly when given only raw visual inputs, across both the five core perceptual dimensions and the two high-level reasoning tasks. Taking Sudoku reasoning as an example, this task requires recognizing digit values and their positions within a grid, and inferring a globally consistent solution. Under the Visual setting, both GPT-4o and Qwen2.5VL-72B fail to handle such constraint-based reasoning, while in the perception setting, accuracy increases from near-zero to 15% and 65%, respectively. This improvement primarily stems from structured perceptual descriptions compensating for the MLLMs’ limited perceptual abstraction capabilities. The difficulty stems from the nature of Sudoku images: small digits, dense layout, and absent semantic cues make them hard for MLLMs to interpret, though humans process such structure effortlessly through rapid visual scanning.
Table 4: Performance Impact of Perceptual Descriptions vs. Visual Inputs in MLLMs. Where ”V” denotes inputting abstract graphics directly, while “P” refers to replacing graphics with structured perceptual descriptions.
Model Num Styl Attr Posit Spat Sudo Rav Avg
GPT-4o(V) 35.0 32.0 38.0 36.0 32.0 0.0 20.0 27.6
GPT-4o(P) 62.0 (+27.0) 53.0 (+21.0) 80.0 (+42.0) 68.0 (+32.0) 100.0 (+68.0) 15.0 (+15.0) 25.0 (+5.0) 60.1(+32.5)
Qwen2.5VL(V) 41.0 43.0 50.0 32.0 40.0 0.0 10.0 30.9
Qwen2.5VL(P) 73.0(+32.0) 83.0(+40.0) 80.0(+30.0) 79.0(+47.0) 100.0(60.0) 65.0(+65.0) 35.0(+25.0) 73.6(+42.7)
Table 5: Ablation study on dataset setting
Model Num Styl Attr Posit Spat Sudo Rav Other Avg
Baseline 23.4 28.2 28.9 23.1 21.7 0.0 14.0 23.9 20.9
+20K 24.5 (+1.1) 29.8 (+1.6) 31.2 (+2.3) 26.1 (+3.0) 23.8 (+2.1) 19.0 (+19.0) 40.0 (+26.0) 22.0 (-1.9) 26.4 (+5.5)
+40K 25.9 (+2.5) 31.8 (+3.6) 34.7 (+5.8) 29.7 (+6.6) 26.4 (+4.7) 30.0 (+30.0) 66.0 (+52.0) 23.5 (-0.4) 30.9 (+10.0)
+60K 26.7 (+3.3) 33.8 (+5.6) 37.8 (+8.9) 33.5 (+10.4) 28.6 (+6.9) 53.0 (+53.0) 87.0 (+73.0) 24.8 (+0.9) 37.2 (+16.3)
Full data 27.2 (+3.8) 35.7 (+7.5) 41.4 (+12.5) 38.3 (+15.2) 30.3 (+8.6) 60.0 (+60.0) 95.0 (+81.0) 21.3 (-2.6) 43.7 (+22.8)
Ablation Study on Perceptually Annotated Synthetic Data. To evaluate the effectiveness of our synthesized data in enhancing model performance on abstract visual reasoning, we conduct a series of ablation studies with incremental training data. The results are shown in Tab. 5. Increasing the amount of synthetic data with perceptual annotations leads to consistent improvements across all dimensions except for the ”Other” category, demonstrating that perceptually structured supervision provides effective guidance for internalizing abstract visual patterns that are difficult to learn from isolated answer-based supervision alone. This further highlights the importance of modeling intermediate perceptual processes in abstract visual reasoning.
5 Conclusion
In this work, we introduce VisuRiddles, a benchmark that systematically evaluates MLLMs’ AVR abilities across diverse, structurally rich tasks, along with an automated data synthesis framework that generates controllable, perceptually annotated reasoning problems. Experimental results demonstrate that even state-of-the-art MLLMs perform only marginally above random chance on VisuRiddles, exhibiting a substantial gap compared to human performance. Our analysis reveals that their limitations stem from inadequate perceptual abstraction rather than reasoning capacity, and that training with our synthetic data improves performance and interpretability. In sum, VisuRiddles enables fine-grained evaluation of reasoning, and our framework offers scalable data support for future training. More importantly, these findings underscore a central insight: enhancing perceptual abstraction is critical to advancing abstract visual reasoning in MLLMs.

### v3

@@
 Abstract
-Recent strides in multimodal large language models (MLLMs) have significantly advanced their performance in many reasoning tasks.
-However, Abstract Visual Reasoning (AVR) remains a critical challenge, primarily due to limitations in perceiving abstract graphics.
-To tackle this issue, we investigate the bottlenecks in current MLLMs and synthesize training data to improve their abstract visual perception.
-First, we propose VisuRiddles, a benchmark for AVR, featuring tasks meticulously constructed to assess models’ reasoning capacities across five core dimensions and two high-level reasoning categories.
-Second, we introduce the Perceptual Riddle Synthesizer (PRS), an automated framework for generating riddles with fine-grained perceptual descriptions.
-PRS not only generates valuable training data for abstract graphics but also provides fine-grained perceptual description, crucially allowing for supervision over intermediate reasoning stages and thereby improving both training efficacy and model interpretability.
-Our extensive experimental results on VisuRiddles empirically validate that fine-grained visual perception is the principal bottleneck and our synthesis framework markedly enhances the performance of contemporary MLLMs on these challenging tasks.
+Recent strides in multimodal large language models (MLLMs) have demonstrated significant progress in many reasoning tasks, but they still fail in Abstract Visual Reasoning (AVR) tasks.
+Our experimental findings indicate that the core bottleneck lies not only in the reasoning capabilities of MLLMs but more critically in their absence of fine-grained perception.
+To address this issue, we present VisuRiddles, a dedicated resource for AVR research.
+It consists of (i) a benchmark, collected from real-world data, for the systematic evaluation of MLLMs’ AVR capabilities, and (ii) a synthesizer, which automatically generates AVR instances enriched with perceptual descriptions and reasoning chains, enabling supervised training and deeper investigation.
+Building on VisuRiddles, we propose a two-stage training paradigm that progressively enhances perceptual ability and strengthens reasoning, producing the Perception-Augmented Visual Reasoner (PAVR).
+Experiments demonstrate that PAVR unifies perception and reasoning, substantially outperforming both open-source and commercial MLLMs, thereby underscoring fine-grained perception as the primary bottleneck in AVR.
 Our code and dataset will be released at https://github.com/yh-hust/VisuRiddles
@@
 (a) Most advanced MLLMs achieve limited accuracy on VisuRiddles, often close to random choice and far below human performance.
-(b) Model responses to abstract graphics and their perceptual descriptions show that once equipped with perceptual understanding, MLLMs can succeed in AVR task.
+(b) Model responses to abstract graphics and their perceptual descriptions show that once equipped with perceptual capability, MLLMs can succeed in AVR task.
 1 Introduction
-Visual reasoning is a core component of human intelligence [1].
-Enhancing this capacity in Multimodal Large Language Models (MLLMs) is key to bridging the gap to human-level performance [10, 11].
-Recent advancements [4, 5, 6, 75, 9, 8, 7] in MLLMs have made significant progress in various general visual tasks, but still struggle with Abstract Visual Reasoning (AVR).
+Reasoning is a core component of human intelligence [1].
+Enhancing this capability in Multimodal Large Language Models (MLLMs) is key to bridging the gap to human-level performance [10, 11].
+Recent advancements [4, 5, 6, 75, 9, 8] in MLLMs have made significant progress in various general visual tasks, but still struggle with Abstract Visual Reasoning (AVR) tasks [84].
 As shown in Fig.
@@
 While such AVR tasks are relatively easy for humans, even the most advanced MLLMs struggle to solve them.
-It seems unexpected that models with strong performance in general understanding and reasoning tasks still fall short in handling such reasoning challenges.
-In fact, the core issue lies not in reasoning but in the fine-grained perception of complex visual structures.
-Unlike humans, who naturally process fine-grained visual structural cues, MLLMs often lack this perceptual ability needed for high-level abstract reasoning.
-As shown in Fig.
-1(b), once the abstract graphics is reformulated into fine-grained perceptual descriptions, the model can produce accurate responses.
-This highlights that enhancing fine-grained perceptual capacity is crucial to unlocking MLLMs’ potential in AVR.
-The significance of AVR tasks derives not only from their inherent complexity but also from their role in revealing the fundamental nature of human perception and cognition [62].
-Therefore, abstract graphics serves as a prototypical and challenging paradigm for evaluating the reasoning capabilities of MLLMs [2, 3].
-Cognitive research suggests that AVR requires the integration of multiple visual dimensions within abstract graphics, including numerosity, attributes, style, position, and spatial relations, which correspond respectively to sensitivity to quantity, object features, stylistic variation, relative placement, and spatial arrangement [63, 64, 65, 66].
-Visual abstraction abilities across these five dimensions form the foundation of visual pattern understanding and serve as indispensable cognitive supports for high-level visual logical reasoning.
-A deficiency in perceiving the visual information of these dimensions can significantly limit a MLLM’s performance on AVR tasks.
-Although solving abstract graphics problems involves multiple cognitive dimensions, benchmarks explicitly designed to evaluate them along these dimensions remain limited.
-To address this gap, we introduce VisuRiddles, a benchmark for evaluating AVR based on the five core dimensions and two high-level reasoning categories.
-VisuRiddles enables a more comprehensive assessment of abstract visual reasoning.
-However, most existing datasets offer only question–answer pairs, lacking the perceptual process supervision needed to bridge visual input and reasoning.
-To tackle this challenge, we further introduce an automated data synthesis framework that generates aligned samples with fine-grained perceptual descriptions, enabling granular supervision across the full perception-to-reasoning pipeline for more systematic learning and improved generalization.
+It unexpected that models with strong performance in general understanding and reasoning tasks still fall short in handling such reasoning challenges.
+In fact, the challenges of AVR primarily stem from two aspects: fine-grained perception and reasoning.
+Compared to the recent advances in enhancing reasoning capabilities, the ability to perceive subtle visual structures, such as position, style, and attribute, has been largely overlooked in existing studies [80, 79].
+Unlike humans, which naturally process such fine-grained visual cues, MLLMs often lack the perceptual ability required for abstract graphics understanding [85, 2].
+As illustrated in Fig.
+1(b), once abstract graphics are reformulated into structured perceptual descriptions, the model is able to generate accurate responses.
+This highlights that strengthening perceptual capability is a crucial factor for advancing AVR.
+Therefore, genuine progress in enabling MLLMs to master AVR requires addressing both perception and reasoning, with special emphasis on enhancing the often-overlooked perceptual capability.
+To advance the study of AVR tasks, we introduce the VisuRiddles benchmark, which is primarily derived from real riddles and enables objective evaluation of MLLMs’ performance.
+Since existing datasets lack fine-grained perceptual annotations, we further design the VisuRiddles Synthesizer, which automatically generates AVR instances enriched with structured perceptual descriptions.
+Leveraging this synthesized data, we perform Supervised Fine-Tuning (SFT) to enhance the model’s ability to perceive fine-grained visual cues, thereby equipping it with the perceptual foundation necessary for AVR.
+Although SFT substantially improves the perceptual capability of MLLMs, they still face persistent bottlenecks in AVR, including errors in perceptual strategy (e.g., difficulty in determining whether the appropriate interpretive cue lies in the figure’s symmetry axis or in its right-angle structures) and limited reasoning ability on more difficult instances.
+Therefore, we introduce Reinforcement Learning (RL), which leverages the enhanced perceptual capability to guide models toward more reliable perceptual grounding and further strengthen their reasoning ability.
+Building upon these components, we develop the Perception-Augmented Visual Reasoner (PAVR), which unifies enhanced perceptual capability with improved reasoning ability to address the dual challenges of AVR.
+Experimental results show that PAVR significantly outperforms advanced commercial models on AVR tasks, highlighting a systematic solution to AVR tasks.
 Our contributions can be summarized as follows:
-  • We introduce VisuRiddles, a comprehensive benchmark for AVR, enabling multi-dimensional evaluation of MLLMs across diverse reasoning dimensions.
-  • We present the Perceptual Riddle Synthesizer, an automated framework designed to address the lack of intermediate supervision by producing structured AVR riddles enriched with fine-grained perceptual descriptions.
-  • We conduct extensive experiments to demonstrate our key findings, including the reasoning limitations of current MLLMs, the effectiveness of our synthesized riddles, and the effective approach to addressing abstract visual problems.
+(1) We introduce VisuRiddles, which comprises a benchmark derived from real riddles for objectively evaluating MLLMs on AVR tasks, and a synthesizer that automatically generates AVR instances enriched with structured perceptual descriptions.
+(2) We leverage synthesized data for SFT to enhance fine-grained perceptual capability for AVR tasks, and further adopt RL to improve reasoning ability and stabilize perceptual strategy selection, resulting in the Perception-Augmented Visual Reasoner.
+(3) We conduct extensive experiments to demonstrate our key findings, including the reasoning limitations of current MLLMs, the effectiveness of our synthesized riddles, and the effective approach to addressing abstract visual problems.
 2 Related Work
-2.1 Multimodal Benchmarks.
+2.1 Multimodal Benchmarks
 High-quality benchmarks are essential for driving progress in MLLMs.
-Earlier evaluation effort predominantly centered on basic visual comprehension tasks, leading to the development of extraction-oriented benchmarks [12, 13, 14, 15, 76, 76, 77, 16].
+Earlier evaluation effort predominantly centered on basic visual comprehension tasks, leading to the development of extraction-oriented benchmarks [12, 13, 14].
 These benchmarks primarily assess a model’s ability to identify and align explicit visual elements.
-However, they depend heavily on surface-level cues, offering little assessment of an MLLM’s capacity for abstraction and logical reasoning.
-Therefore, researchers have expanded their focus on more holistic evaluation frameworks, proposing general-purpose benchmarks [17, 18, 19, 20, 78].
-These benchmarks span multiple tasks, including visual understanding, mathematical reasoning, and knowledge-based QA, and are designed to assess the generalist capabilities of MLLMs.
-Although general-purpose benchmarks span a wide range of tasks, they tend to emphasize knowledge-based evaluation over visual logical reasoning.
+However, they depend heavily on surface-level cues, offering little assessment of an MLLM’s capability for abstraction and logical reasoning.
+Therefore, general-purpose benchmarks [17, 19, 86, 20, 18] are widely focused on.
+These benchmarks span multiple tasks, such as visual understanding and mathematical reasoning, and are designed to assess the generalist capabilities of MLLMs.
+However, they tend to emphasize knowledge-based evaluation over visual logical reasoning.
 Moreover, due to their strong semantic dominance, MLLMs may directly rely on their inherent knowledge, making it difficult to disentangle and evaluate models’ true visual reasoning capabilities.
-To address these limitations, certain studies have shifted toward evaluating visual logical reasoning, leading to logic-oriented benchmarks such as PGM [24], RAVEN [21], I-RAVEN [22], RAVEN-FAIR [23], CVR [25], Raven IQ [26], ConceptARC [27], RPMs [28], MaRs-VQA [2] MathVista [29], GeoEval [30], MathVision[31], MathVerse [32], CMMath [33], OCR-Reasoning [78] and MV-MATH [34].
+Therefore, some studies have shifted toward evaluating visual logical reasoning, leading to logic-oriented benchmarks such as RAVEN [21], I-RAVEN [22], RAVEN-FAIR [23], CVR [25], RPMs [28], MaRs-VQA [2], MathVerse [32], MARVEL [79], PuzzleVQA [83], VisCogBench [2], VisuLogic [80], VisualPuzzle [81], VisualSphinx [82] and MV-MATH [34].
 These benchmarks use structured diagrams, mathematical forms, and spatial layouts to assess analogical reasoning, compositionality, and pattern induction, signaling a move toward higher-level AVR.
 However, these benchmarks still exhibit notable limitations, particularly in their partial reliance on external knowledge and lack breadth in reasoning coverage.
-2.2 Multimodal Large Language Models for Visual Reasoning.
+2.2 Multimodal Large Language Models for Visual Reasoning
 Advanced commercial LLMs, including the GPT [35], Gemini [36], and Claude [37] series, have exhibited impressive capabilities in multimodal understanding and generation.
-These advances have stimulated increasing interest and systematic investigation into MLLMs from both academia and industry.
-Early representative general-purpose MLLMs, such as LLaVA [38], Instruct-BLIP [39], Qwen-VL [40], Intern-VL [41], and mPLUG-Owl [42], perform well on general multimodal tasks but exhibit limitations in fine-grained visual perception.
-Subsequently, DeepSeek-VL [43], Monkey [8], InternVL 1.5 [44], TextMonkey [45], InternLM-XComposer2-4KHD [46], Qwen2-VL [47], and mPLUG-Owl2 [48] have introduced high-resolution visual perception mechanisms, establishing a more robust foundation for fine-grained visual understanding and reasoning.
+These advances have spurred systematic research on MLLMs in both academia and industry.
+Early representative general-purpose MLLMs, such as LLaVA [38], Instruct-BLIP [39], Qwen-VL [40], and Intern-VL [41], perform well on general multimodal tasks but exhibit limitations in fine-grained visual perception.
+Subsequently, DeepSeek-VL [43], Monkey [8], TextMonkey [45], InternLM-XComposer2-4KHD [46], Qwen2-VL [47] have introduced high-resolution visual perception mechanisms, establishing a more robust foundation for fine-grained visual understanding and reasoning.
 However, such improvements primarily expand the perceptual scope of MLLMs without substantially enhancing their reasoning capabilities.
-Therefore, inference-time scaling methods including LLaVA-CoT[49], BBA [50], R-CoT [51], and RedStar [52] have introduced and innovated upon inference-time scaling techniques such as chain-of-thought (CoT), significantly enhancing the visual reasoning capabilities of MLLMs.
+Therefore, inference-time scaling methods including MMVP [85], LLaVA-CoT [49], BBA [50], R-CoT [51], and RedStar [52] have introduced and advanced inference-time scaling techniques such as chain-of-thought (CoT), significantly enhancing the visual reasoning capabilities of MLLMs.
 More recently, following the advances of Kimi K1.5 [53] and DeepSeek-R1 [54], a series of studies, such as R1-OneVision [3], LMM-R1 [55], MM-EUREKA [56], R1-V [57], Visual-RFT [58], VisualPRM [59], OThink-MR1 [60], and VLM-R1 [61], have employed reinforcement learning to develop adaptive reasoning policies and improve generalization in complex visual reasoning tasks.
-Figure 2: Representative examples from VisuRiddles. The benchmark includes eight reasoning categories, designed to comprehensively evaluate diverse reasoning capabilities of MLLMs.
-3 Method
-In this section, we first describe the construction process of VisuRiddles, following a structured pipeline consisting of data collection, cleaning, and consolidation.
-We then summarize key benchmark statistics—including question and image counts, as well as answer distribution in Sec.
-3.1.
-Subsequently, we introduce the Riddles Synthesis Framework, an automated approach for generating structurally diverse and perceptually annotated samples tailored for AVR in Sec.
-3.2.
+3 VisuRiddles
+In this section, we introduce VisuRiddles, a innovative and comprehensive resource for AVR research.
+It consists of the VisuRiddles Benchmark (Sec.
+3.1), built from real-world problems to objectively evaluate MLLMs’ AVR performance, and the VisuRiddles Synthesizer (Sec.
+3.2), designed to generate diverse perceptually annotated AVR instances for model training.
 3.1 VisuRiddles Benchmark
-Existing benchmarks provide limited coverage of the core reasoning dimensions required for AVR, To address this, VisuRiddles explicitly evaluates models across five key dimensions: a) Numerosity assesses the model’s ability to perceive and reason about quantity and distribution; b) Attribute evaluates the understanding of intrinsic visual features that determine structural semantics; c) Style tests the capacity to identify and generalize transformation-based visual patterns; d) Position reflects the ability to reason over the relative positions and layout of visual elements; e) Spatiality examines the understanding of three-dimensional structures, shape variations, and spatial transformations in abstract scenes.
+Unlike RAVEN [21] and PuzzleVQA [83], which focus on general visual reasoning tasks, we target the most challenging part of AVR tasks.
+Thus, the VisuRiddles benchmark has been constructed to evaluate MLLMs.
+It explicitly evaluates models across five key dimensions: a) Numerosity assesses the model’s ability to perceive and reason about quantity and distribution; b) Attribute evaluates the understanding of intrinsic visual features that determine structural semantics; c) Style tests the capability to identify and generalize transformation-based visual patterns; d) Position reflects the ability to reason over the relative positions and layout of visual elements; e) Spatiality examines the understanding of three-dimensional structures, shape variations, and spatial transformations in abstract graphics.
 To extend the evaluation beyond low-level perceptual reasoning and address limitations in task diversity, structural complexity, and the inherent uncertainty introduced by single-choice formats, we additionally incorporate two high-level reasoning tasks that require analogical abstract reasoning and consistency-based logical reasoning, as represented by RAVEN and Sudoku, respectively.
@@
 2.
+We construct the VisuRiddles benchmark through a structured pipeline comprising Collection, Cleaning, and Consolidation.
+We provide the construction details in Appendix A.
+Figure 2: Representative examples from VisuRiddles. The benchmark includes eight reasoning categories, designed to comprehensively evaluate diverse reasoning capabilities of MLLMs.
+We initially collected 1,275 AVR problems, each accompanied by expert-level analysis and answer.
+After careful filtering and quality control, we construct VisuRiddles with 800 samples in the basic categories and 200 samples in the high-level categories.
+Key statistics of VisuRiddles are summarized in Tab.
+1.
+For the basic AVR tasks, each question is formulated as a single-choice problem, with answer option distributions of A (25.75%), B (24.75%), C (24.375%), and D (25.125%).
+In the high-level categories, RAVEN Reasoning and Sudoku Reasoning each account for 10% of the total data, and require models to generate exact symbolic outputs to be considered correct.
+Table 1: The statistics of VisuRiddles, including the number questions, average question length (in tokens), number of associated images, and the corresponding answer format.
+Type Questions Number Question length Answer Formats
+Sudoku 100 48.3 Constraint-based grid
+RAVEN 100 70.7 Single-choice
+Numerical 250 26.3 Single-choice
+Stylistic 117 26.3 Single-choice
+Attribute 97 27.0 Single-choice
+Positional 111 25.5 Single-choice
+Spatial 156 27.2 Single-choice
+Other 69 30.3 Single-choice
+Total 1000 32.5 -
+Figure 3: Overview of the VisuRiddles Synthesizer.
+(a) A unified pipeline for generating abstract graphics with fine-grained perceptual descriptions.
+(b) Visualization of synthesized riddles based on positional rule and stylistic rule.
+3.2 VisuRiddles Synthesizer
+Recent MLLMs demonstrate strong general reasoning, but their limited performance on AVR is largely attributed to insufficient capability of fine-grained perception for abstract graphics.
+Most existing datasets lack intermediate perceptual description, providing only question-answer pairs [80, 79, 81].
+This prevents explicit modeling of the perception-to-reasoning process, leading to black-box inference, weak inductive capability, and poor generalization.
+More critically, existing commercial models struggle to annotate perceptual processes effectively, and manual labeling requires significant expert effort, both of which further exacerbate the issue.
+To address this, we introduce VisuRiddles Synthesizer, a Riddles Synthesis Framework to provides abstract graphics with aligned fine-grained perceptual description.
+Figure 3 illustrates the pipeline of our VisuRiddles Synthesizer for AVR task.
+The VisuRiddles Synthesizer pipeline comprises two stages: Riddles Construction and API Labeling.
+The former generates abstract visual instances with aligned perceptual descriptions, while the latter generates reasoning chains based on these descriptions.
+Details and representative examples are provided in Appendix B.2 and Appendix B.3.
+The VisuRiddles Synthesizer generates data that spans five core perceptual reasoning types and two high-level reasoning tasks.
+The detailed configurations of VisuRiddles Synthesizer are provided in Table 5.
+The VisuRiddles Synthesizer not only provides riddles’ visual instances but also structured perceptual descriptions, thereby enhancing the model’s fine-grained perceptual ability for AVR tasks.
+It is noteworthy that the riddles instances generated by the VisuRiddles Synthesizer focus on improving fine-grained perception, rather than reasoning.
+Therefore, their reasoning difficulty is deliberately kept lower than that of real-world riddles.
+4 Perception-Augmented Visual Reasoner
+To tackle the challenges of AVR, we introduce the Perception-Augmented Visual Reasoner (PAVR).
+As illustrated in Fig.
+4, PAVR is developed through a two-stage training paradigm.
+In the first stage, SFT equips the model with the ability to capture fine-grained visual cues in abstract graphics.
+Building upon this, the second stage employs RL to effectively integrate these visual cues into stable reasoning strategies for solving complex tasks, resulting in PAVR.
+These two stages are complementary: SFT enables the model to see clearly, while RL guides it to reason reliably.
+Figure 4: Overview of PAVR.
+(i) Baseline: Incorrect perception leads to incorrect results.
+(ii) PAVR-SFT, which is trained on synthesized data to enhance fine-grained perceptual ability, can accurately understand visual content in riddles but fails in pattern identification.
+(iii) PAVR, which builds upon PAVR-SFT with reinforcement learning, can effectively recognize patterns and derive the correct answer.
+Perceptual Augment via Supervised Fine-Tuning.
+The lack of fine-grained perceptual ability remains a major bottleneck for MLLMs on AVR tasks, yet existing real-world datasets typically provide only single-choice answer and rarely provide explicit perceptual annotations to supply sufficient supervision.
+To address this limitation, we employ the VisuRiddles Synthesizer described in Sec.
+3.2 to generate AVR instances across seven categories, each enriched with structured perceptual descriptions.
+These synthesized examples serve as training data for SFT, enhancing the model’s perceptual ability and enabling it to capture fine-grained cues in abstract graphics.
+This establishes a reliable foundation for subsequent reasoning optimization through reinforcement learning.
+Reasoning Optimization via Reinforcement Learning.
+Although SFT on synthesized data improves fine-grained perceptual ability, it still faces limitations in AVR, including unstable perceptual strategy selection (e.g., axis vs.
+angles) and insufficient reasoning on complex tasks.
+Therefore, we apply Group Relative Policy Optimization (GRPO) [54] to endow the model with more reliable perceptual grounding and stronger reasoning ability.
+The reward design includes: (i) Answer Reward, giving 1 for correct answers and 0 otherwise, and (ii) Format Reward, incentivizing outputs that match the required template: <think> … </think> <answer> … </answer>.
+The training data for GRPO is likewise generated by the VisuRiddles Synthesizer.
+This process enables the model to achieve more reliable perceptual grounding and stronger reasoning capabilities.
+5 Experiment
+5.1 Experimental Setup
+We use Qwen2.5-VL-7B [4] as the baseline model to develop PAVR.
+In the supervised fine-tuning stage, the model is trained for 20 epochs using 20K synthesized AVR instances generated by the VisuRiddles Synthesizer.
+Subsequently, in the GRPO stage, optimization is performed using 4K synthesized instances over 40 epochs.
+All experiments are conducted on a setup comprising 8 NVIDIA A800 80G GPUs.
+Evaluation settings are summarized in Appendix C.
+Table 2: Evaluation result on VisuRiddles benchmark.
+The superscript number indicates the number of answer options (e.g., “4” means a 4-choice question with one correct option; * represents a vast combinatorial solution space).
+The best results are marked bold and the second results are underlined.
+Model Param Num4 Styl4 Attr4 Posit4 Spat4 Sudo* Rav8 Other4 Avg
+Human - 61.3 60.9 67.5 67.9 58.8 - - 61.9 -
+Open-Source MLLMs
+Minicpm-V-2.6 8B 21.6 26.5 23.7 31.5 23.1 0.0 10.0 24.6 20.6
+InternVL2.5-8B 8B 21.2 13.7 30.9 23.4 23.1 0.0 0.0 29.0 18.1
+InternVL2.5-8B-MPO 8B 22.0 23.9 27.8 25.2 21.8 0.0 3.0 27.5 19.4
+Deepseekvl2 27B 21.2 27.4 18.6 18.0 17.3 0.0 15.0 13.0 17.4
+InternVL2.5-38B 38B 22.0 19.7 29.9 26.1 29.5 1.0 19.0 30.4 22.3
+InternVL2.5-38B-MPO 38B 26.0 20.5 19.6 27.0 27.6 0.0 18.0 31.9 22.1
+Qwen2.5VL-32B 32B 23.2 20.5 30.9 27.0 29.5 0.0 40.0 24.6 24.5
+InternVL2.5-78B 78B 26.0 25.6 27.8 29.7 23.7 0.0 16.0 27.5 22.7
+InternVL2.5-78B-MPO 78B 25.6 30.8 24.7 27.9 26.3 0.0 12.0 20.3 22.2
+InternVL2.5-78B-MPO(cot) 78B 23.2 23.1 24.7 28.8 22.4 0.0 11.0 23.2 20.3
+Qwen2.5VL-72B 72B 23.6 23.1 19.6 30.2 26.9 0.0 62.0 23.9 25.9
+Qwen2.5VL-72B(cot) 72B 27.2 26.5 24.7 19.8 28.2 0.0 55.0 23.2 26.0
+Qwen3-VL-235B-A22B-Instruct 235B 31.2 28.2 36.1 34.2 26.3 29.0 53.0 27.5 32.6
+Qwen3-VL-235B-A22B-Thinking 235B 31.2 29.9 44.3 33.3 30.1 33.0 49.0 39.1 34.9
+Dots.vlm1 671B 30.4 31.6 38.1 38.7 35.3 1.0 20.0 33.3 29.2
+Commercial products
+GPT-4o - 23.2 21.4 29.9 27.9 28.2 0.0 14.0 24.6 21.8
+GPT-4o(cot) - 22.4 23.9 32.0 29.7 30.1 0.0 22.0 29.0 23.7
+o3 - 28.4 40.2 41.2 27.0 21.8 0.0 25.0 33.3 27.0
+claude-3-7-sonnet - 24.0 27.4 27.8 29.7 28.2 4.0 43.0 31.9 26.5
+claude-3-7-sonne(cot) - 24.4 36.8 27.8 29.7 25.6 3.0 36.0 27.5 26.2
+claude-3-7-sonnet-thinking - 24.0 24.8 33.0 23.4 25.6 3.0 24.0 26.1 23.2
+Gemini-2.5-flash-thinking - 21.2 30.8 27.8 21.6 16.7 17.0 16.0 23.2 21.5
+Gemini2.5-pro - 31.6 31.6 48.5 26.1 30.1 39.0 30.0 44.9 33.9
+GPT-5 - 30.8 30.8 38.1 32.4 30.8 2.0 29.0 31.9 28.7
+Ours
+Baseline(Qwen2.5VL-7B) 7B 24.4 28.2 23.7 22.5 25.0 0.0 48.0 24.6 24.6
+PAVR-SFT 7B 31.2 31.6 44.3 31.5 45.5 43.0 61.0 39.1 39.5
+PAVR 7B 39.6 39.3 50.5 39.6 51.9 46.0 65.0 55.1 46.8
+5.2 Main Results
+Table 2 presents performance of various MLLMs on VisuRiddles benchmark. We summarize the key observations as follows:
+Performance of Open-Source MLLMs.
+Most open-source MLLMs demonstrate limited performance on VisuRiddles, with accuracies across core categories often close to random choice and consistently failing on the two high-level reasoning tasks, which reveals a substantial gap between current model capabilities and human performance in AVR.
+Notably, the newly released Qwen3-VL-235B-A22B-Instruct and Qwen3-VL-235B-A22B-Thinking exhibit markedly stronger results, though such improvements remain exceptions among open-source models.
+Performance of Commercial Models with Thinking Mode.
+Several commercial MLLMs with ‘thinking’ mode achieve comparatively improved performance on VisuRiddles.
+Their advantage is particularly notable in structured tasks like Sudoku and Raven.
+These results indicate that AVR can benefit from advanced reasoning capabilities.
+However, the overall performance remains limited.
+Effectiveness of Model Scaling and CoT Prompting.
+Strategies such as increasing model size and applying CoT prompting, though commonly used to enhance reasoning, are ineffective for AVR tasks.
+Models with larger parameter scales do not consistently outperform smaller ones, and CoT yields only limited improvements on select tasks.
+These findings indicate that scaling in model parameter or inference time is insufficient to address the core challenges of AVR.
+Effectiveness of PAVR.
+PAVR significantly outperforms other MLLMs across all categories of VisuRiddles, especially in high-level reasoning tasks, demonstrating the model’s ability to handle complex visual patterns.
+Notably, compared to the baseline, the significant improvement achieved by training on data with perceptual descriptions through SFT highlights the critical role of fine-grained perceptual ability in AVR tasks.
+On the other hand, the application of RL further enhances performance across various dimensions on VisuRiddles, with the improvements primarily stemming from enhanced reasoning capabilities through strategic guidance.
+We also provide the results of PAVR on the VisuLogic [80] benchmark in  D. The experimental results further support our conclusions.
+5.3 Visualization Results
+Figure 5: Case study comparing different reasoning strategies on a VisuRiddles example.
+(a) QVQ-72B reflects a flawed loop under the "thinking" mode.
+(b) Qwen2.5-VL-72B with CoT prompting exhibits incorrect perceptual understanding.
+(c) PAVR exhibits accurate perception and coherent reasoning, ultimately arriving at the correct answer.
+To further analyze the reason that MLLMs underperform on AVR task, we examine two widely adopted reasoning enhancement strategies: inference-time scaling and CoT prompting.
+We conduct experiments on QVQ-72B-Preview (equipped with a "think" mode), Qwen2.5-VL-72B (with CoT prompting), and PAVR.
+Fig.
+5 visualizes a representative case of a broadly observed phenomenon uncovered during our experiments.
+The results reveal clear differences in the reasoning behaviors across these models.
+The inference-time scaling approach, which leverages "think"-style prompting without any perceptual grounding, frequently leads to verbose yet logically inconsistent reasoning, often ending in incorrect answers.
+The CoT prompting strategy often suffers from perceptual errors, which cascade into flawed reasoning and ultimately lead to incorrect answers.
+In contrast, by grounding reasoning in fine-grained perceptual capability, PAVR correctly interprets the riddle and identifies the underlying pattern, ultimately producing the correct answer.
+Notably, while PAVR occasionally exhibits erroneous perception, we observe a ‘rethink’ phenomenon, where the model re-examines and corrects them, realigning the reasoning trajectory to reach the correct solution.
+These results underscore the importance of perception in AVR.
+Inference-time scaling and CoT prompting alone cannot compensate for the lack of perceptual grounding, which often leads to incorrect or inconsistent reasoning.
+Conversely, the incorporation of fine-grained perceptual capability enables the model to capture subtle visual cues and sustain reliable performance on AVR tasks.
+5.4 Ablation Study
+Ablation Study on Bottlenecks in AVR.
+To investigate whether the performance gap in AVR stems from perceptual limitations rather than reasoning ability, we conduct an ablation study using our perception-annotated synthetic data.
+We evaluate two representative MLLMs, GPT-4o (closed-source) and Qwen2.5VL-72B (open-source), on a subset of our synthetic dataset under two input settings: (i) raw abstract graphics (V), and (ii) structured perceptual descriptions (P).
+The experimental results in Tab.
+3 show that both models perform poorly when given only raw visual inputs.
+Taking Sudoku reasoning as an example, this task requires recognizing digit values and their positions within a grid, and inferring a globally consistent solution.
+Under the Visual setting, both GPT-4o and Qwen2.5VL-72B fail to handle such constraint-based reasoning, while in the perceptual descriptions setting, accuracy increases from near-zero to 15% and 65%, respectively.
+This improvement primarily stems from structured perceptual descriptions compensating for the MLLMs’ limited perceptual abstraction capabilities.
+The difficulty stems from the nature of Sudoku images: small digits, dense layout, and absent semantic cues make them hard for MLLMs to interpret, though humans process such structure effortlessly through rapid visual scanning.
+Table 3: Impact of Descriptions vs. Visual Inputs in MLLMs. Where "V" denotes inputting abstract graphics directly, while “P” refers to replacing graphics with structured perceptual descriptions.
+Model Num. Styl. Attr. Posit. Spat. Sudo. Rav. Avg
+GPT-4o(V) 35.0 32.0 38.0 36.0 32.0 0.0 20.0 27.6
+GPT-4o(P) 62.0 (+27.0) 53.0 (+21.0) 80.0 (+42.0) 68.0 (+32.0) 100.0 (+68.0) 15.0 (+15.0) 25.0 (+5.0) 60.1(+32.5)
+Qwen2.5VL(V) 41.0 43.0 50.0 32.0 40.0 0.0 10.0 30.9
+Qwen2.5VL(P) 73.0(+32.0) 83.0(+40.0) 80.0(+30.0) 79.0(+47.0) 100.0(60.0) 65.0(+65.0) 35.0(+25.0) 73.6(+42.7)
+Ablation Study on Reasoning Enhancements in AVR.
+To examine the contribution of reasoning-oriented enhancements, we conduct an ablation study on three components: Caption, CoT, and GRPO, as shown in Table 4.
+We can draw the following conclusions form the experiments: (i) perception serves as the essential foundation, as reasoning enhancements alone contribute only marginal improvements; (ii) relying solely on structured perceptual descriptions for SFT yields limited gains and weak generalization, whereas augmenting SFT with CoT annotations generated by MLLMs alleviates this issue, endowing the model with fine-grained perceptual ability and partial reasoning capability; and (iii) Building on a perception-grounded model, the incorporation of GRPO further unifies perception and reasoning, ultimately delivering substantial performance improvements.
+Table 4: Ablation study on Reasoning Enhancements.
+Where Caption denotes training with fine-grained perceptual descriptions, CoT denotes augmenting training with chain-of-thought reasoning traces that incorporate structured perceptual descriptions generated by MLLMs, and GRPO refers to applying Group Relative Policy Optimization to strength model’s reasoning capability.
+Model SFT RL Avg
+Baseline (Qwen2.5-VL) 24.6
+Baseline + Caption 33.3 (+8.7)
+Baseline + GRPO 29.4(+4.8)
+Baseline + CoT (PAVR-SFT) 39.5(+14.9)
+Baseline + CoT + GRPO (PAVR) 46.8 (+22.2)
+6 Conclusion
+In this work, we observe that the primary bottleneck of MLLMs in AVR lies in their overlooked fine-grained perceptual capability.
+To address this issue, we introduce VisuRiddles, a unified resource that comprises a benchmark for objectively evaluating MLLMs on AVR tasks and a synthesizer for generating training instances with structured perceptual descriptions.
+Building on this resource, we propose PAVR, which unifies perception and reasoning through supervised fine-tuning for perceptual enhancement and GRPO-based optimization for reasoning improvement for AVR tasks.
+Experimental results demonstrate that current MLLMs perform poorly on AVR tasks.
+Moreover, scaling model parameters, CoT prompting, and inference-time scaling fail to effectively address this challenge.
+These findings highlight that fine-grained perception constitutes the fundamental bottleneck in AVR, while reasoning enhancements applied based on a perception-grounded model can further improve performance.
+In short, our work not only offers a systematic resource for both evaluation and training, but also provides a practical framework for addressing the AVR challenge.
+Limitation: Due to limitations in time and human labor, the scale of the resources (e.g., rule design, icon library) used for data synthesis is restricted, which somewhat limits the diversity and richness of the synthetized data.
+In future work, we plan to expand the resource pool to a larger scale with greater complexity, further enriching the synthesized riddles.
+APPENDIX
+Appendix A Construction Details of VisuRiddles Benchmark
 We construct the VisuRiddles benchmark through a structured pipeline comprising Collection, Cleaning, and Consolidation:
@@
 We manually collect a diverse set of visual reasoning problems from Chinese National Civil Service Examination, which form the five core categories of the benchmark.
-RAVEN and Sudoku samples are generated following the method1111https://github.com/WellyZhang/RAVEN1 and the protocols2222https://gitee.com/mxx11/sudoku2.
+RAVEN and Sudoku samples are generated following the method [^1] https://github.com/WellyZhang/RAVEN and the protocols [^2] https://gitee.com/mxx11/sudoku.
 Cleaning.
@@
 All data were standardized into a unified JSON format, supporting structured access to image content, metadata, and reasoning annotations.
-Following the structured pipeline, we initially collected 1,275 abstract visual reasoning problems, each accompanied by expert-level analysis and answer.
-After careful filtering and quality control, we construct VisuRiddles with 770ok samples in the basic categories and 200 samples in the high-level categories.
-Key statistics of VisuRiddles are summarized in Tab.
-2.
-For the basic AVR tasks, each question is formulated as a single-choice problem, with answer option distributions of A (25.75%), B (24.75%), C (24.375%), and D (25.125%).
-In the high-level categories, RAVEN Reasoning and Sudoku Reasoning each account for 10% of the total data, and require models to generate exact symbolic outputs to be considered correct.
-Table 1: The statistics of VisuRiddles, including the number questions, average question length (in tokens), number of associated images, and the corresponding answer format.
-Type Questions Number Question length Image Number Answer Formats
-Sudoku Reasoning 100 48.3 100 Constraint-based grid
-RAVEN Reasoning 100 70.7 200 Pattern match index
-Numerical Reasoning 250 26.3 250 Single-choice
-Stylistic Reasoning 117 26.3 117 Single-choice
-Attribute Reasoning 97 27.0 97 Single-choice
-Positional Reasoning 111 25.5 111 Single-choice
-Spatial Reasoning 156 27.2 156 Single-choice
-Other 69 30.3 69 Single-choice
-Total 1000 32.5 1100
-3.2 Perceptual Riddle Synthesizer
-Recent MLLMs demonstrate strong general reasoning, but their limited performance on AVR is largely attributed to insufficient capacity of fine-grained perception.
-Most existing datasets lack intermediate perceptual description, providing only question-answer pairs.
-This prevents explicit modeling of the perception-to-reasoning process, leading to black-box inference, weak inductive capacity, and poor generalization.
-More critically, annotating such perceptual processes requires effort of expert-level staff members, further compounding this problem.
-To address this, we introduce Perceptual Riddle Synthesizer (PRS), a Riddles Synthesis Framework to provides abstract graphics with aligned perceptual description.
-Figure 3 illustrates the pipeline of our PRS framework for AVR task. The pipeline consists of three main stages:
-Rule Selection.
+Appendix B Details of VisuRiddles Synthesizer
+B.1 Configuration of the VisuRiddles Synthesizer
+Table 5: Configuration details of the VisuRiddles Synthesizer across seven reasoning categories.
+Rule Sub-rule Layout Template Icon
+Numerical Line, Curve, Angle, Cart, Space, Parts 1 1 5k
+Stylistic AND, OR, XOR, XNOR 3 5 5k
+Attribute Element, Group 2 1 5k
+Positional Translate, Rotate, Flip 4 7 5k
+Spatial Unfolding, Three-View, 3D-Reconstruction 7 13 5k
+View-Consistency, Multiple-Views
+Sudoku Sudoku 1 5 9
+RAVEN Raven 1 7 10k
+B.2 Synthesis Pipeline of Riddles Instances
+The synthesis pipeline for generating riddle instances with structured perceptual descriptions consists of two main stages: Instance Construction and API Labeling.
+Instance Construction
+Instance synthesis is responsible for generating the visual instances and corresponding perceptual descriptions. The detailed steps are as follows:
+1. Rule Selection:
 The synthesis process begins with rule selection.
@@
 This decomposition enables modular and interpretable synthesis, allowing each instance to target a specific reasoning skill.
-Element Configuration.
+2. Element Configuration:
 Once a sub-rule is selected, we proceed to define the visual components in a structured manner.
-This includes: (i) choosing a layout template (e.g., $3\times 33 × 3$ matrix, progression row, etc.), (ii) selecting or auto-generating icon elements with specific visual properties (e.g., shape, color, direction), and (iii) assigning a background grid.
+This includes: (i) choosing a layout template (e.g., $3\times 3$ matrix, progression row, etc.), (ii) selecting or auto-generating icon elements with specific visual properties (e.g., shape, color, direction), and (iii) assigning a background grid.
 Elements are configured in a rule-consistent way: for example, a “positional” sub-rule may require translation across rows, while a “stylistic” sub-rule may apply logical operations to shape fill.
 Each configuration ensures the pattern remains visually interpretable and solvable.
-Rendering and Annotating.
+3. Rendering and Annotating:
 Each question instance is constructed by applying parameterized transformations under defined constraints: a single correct answer, consistent application of one rule, and plausible distractors.
@@
 These annotations bridge perception and reasoning, enabling models to explicitly learn visual abstraction rather than relying on black-box pattern matching.
-PRS produces a large-scale dataset spanning five core perceptual reasoning types and two high-level structured reasoning tasks.
-It supports 34 sub-rules across these categories and generates samples using 8 layout and 10k unique graphical icons.
-Each instance is created through rule-guided visual composition and rendered alongside structured perceptual annotations.
-The resulting dataset exhibits high diversity in both visual patterns and reasoning logic, as summarized in Tab.
-2.
-It is worth noting that the synthetic tasks are designed to strengthen perceptual abstraction rather than replicate real-world reasoning difficulty.
-Consequently, their complexity is deliberately kept tractable to focus on providing structured perceptual supervision.
-Figure 3: Overview of the PRS framework.
-(a) A unified pipeline for generating abstract graphics with fine-grained perceptual descriptions across different rule categories.
-(b) Visualization of synthesized riddles based on positional rule and stylistic rule.
-Table 2: Configuration details of PRS across seven reasoning categories.
-Rule Sub-rule Layout Template Icon Sample
-Numerical 5 2 1 5k 10K
-Stylistic 6 3 5 5k 10K
-Attribute 2 2 1 5k 10K
-Positional 8 4 7 5k 10K
-Spatial 5 2 13 5k 10K
-Sudoku 1 1 4 9 10K
-RAVEN 7 5 5 10k 20K
-4 Experiment
-In this section, we present extensive experiments on advanced MLLMs on the challenging VisuRiddles benchmark.
-We first describe the experimental setup for both training and evaluation in Sec.
-4.1.
-Subsequently, we demonstrate a comprehensive evaluation of MLLMs in Sec.
-4.2.
-We further include visualization examples that illustrate model behaviors and failure modes in Sec.
-4.3.
-Finally, we conduct ablation studies to validate the efficacy of our synthesis engine in Sec.
-4.4.
-4.1 Experimental Setup
-Training details.
-We fine-tune the open-source model Qwen2.5-VL-7B using 80k perceptually annotated samples generated by our Riddles Synthesis Framework.
-Training is performed with a learning rate of 5e-6, using the Adam optimizer, a batch size of 16, and 8 NVIDIA A800 GPUs.
-Evaluation Settings.
-We evaluate seventeen MLLMs, including eleven open-source and six advanced proprietary models.
-The open-source group consists of MiniCPM-V 2.6 [7], DeepSeek-VL2 [67], Qwen2.5-VL [68], InternVL2.5 [5] and InternVL2.5-MPO (Mixed Preference Optimization) [69], which is designed to improve reasoning capabilities via reinforcement learning-based optimization.
-The proprietary group includes GPT-4o (20241120) [70], o1 [71], Claude-3-7-Sonnet (20250219) [72], Claude-3-7-Sonnet-Thinking (20250219) [72], Gemini-2.0-Flash-Thinking (20250121) [73], and Gemini-2.5-Pro (20250325) [74].
+API Labeling
+Due to the poor performance of existing commercial models in abstract visual reasoning (AVR) tasks, especially when processing images, these models often suffer from significant hallucination issues and fail to generate correct reasoning annotations.
+Thus, directly feeding images into the model does not produce accurate results.
+To overcome this, we introduce a new step: first, we replace the images with corresponding perceptual descriptions and input these descriptions into a large model.
+This allows the model to generate CoT based on the perceptual descriptions.
+The detailed steps are as follows:
+1. CoT Generation:
+Perceptual descriptions are first extracted from the synthesized visual instances to supply fine-grained visual information.
+These descriptions are then provided to an advanced commercial model (Gemini 2.5-Flash-Think [^3] https://deepmind.google/models/gemini/flash/ in this work), which generates a Chain-of-Thought (CoT) reasoning process and produces a corresponding answer.
+2. Validation:
+The generated answer is compared to the correct answer of the synthesized riddle.
+If the model’s output does not align with the correct answer, the data point is discarded, ensuring that only valid and accurate samples are included in the final dataset.
+B.3 Examples of Synthesized Riddles Instances
+Figure 6: Examples of synthesized riddle instances in Numerical, Positional, and Attribute
+Figure 7: Examples of synthesized riddle instances in Spatial, Stylistic, Sudoku, and RAVEN
+Appendix C Details of Evaluated Models
+We evaluate seventeen MLLMs, including eleven open-source models and six advanced proprietary models, as shown in Table  6.
+Human reference answers are obtained from the Fenbi Educational Platform, a widely used source for civil service exam preparation.
 Notably, GPT-o1, Claude-3-7-Sonnet-Thinking, and Gemini-2.5-Pro are equipped with specialized “thinking” modes aimed at enhancing multi-step reasoning.
@@
 Inference for all models is performed using the vLLM framework with unified decoding configurations.
-Human reference answers are obtained from the Fenbi Educational Platform, a widely used source for civil service exam preparation.
-Table 3: Evaluation of various models on VisuRiddles.
-The superscript number indicates the number of answer options (e.g., “4” denotes a 4-choice question where one option is correct; $\infty$” represents unlimited possible options).
-The best results are marked bold and the second results are underlined.
+Table 6: Overview of evaluated MLLMs.
+Model Name Params Think CoT Provider Version
+Open-Source MLLMs
+Qwen2.5-VL-7B [68] 7B Alibaba -
+Qwen2.5-VL-32B 32B Alibaba -
+Qwen2.5-VL-72B 72B Alibaba -
+Qwen2.5-VL-72B(CoT) 72B Alibaba -
+Qwen3-VL-235B-A22B-Instruct 235B Alibaba -
+Qwen3-VL-235B-A22B-Thinking 235B Alibaba -
+InternVL2.5-8B [5] 8B Shanghai AI Lab -
+InternVL2.5-MPO-8B [69] 8B Shanghai AI Lab -
+InternVL2.5-38B 38B Shanghai AI Lab -
+InternVL2.5-38B-MPO 38B Shanghai AI Lab -
+InternVL2.5-78B 78B Shanghai AI Lab -
+InternVL2.5-78B-MPO 78B Shanghai AI Lab -
+InternVL2.5-78B-MPO(CoT) 78B Shanghai AI Lab -
+DeepSeek-VL2 [67] 32B DeepSeek -
+MiniCPM-V 2.6 [7] 8B ModelBest -
+Dots.vlm1 671B Rednote hi lab -
+Commercial Products
+GPT-4o - OpenAI 2024.11.20
+GPT-4o(CoT) - OpenAI 2024.11.20
+o3 - OpenAI 2025.04.16
+Claude-3.7-Sonnet - Anthropic 2025.02.19
+Claude-3.7-Sonnet-Thinking - Anthropic 2025.02.19
+Gemini-2.0-Flash-Thinking - Google DeepMind 2025.01.21
+Gemini-2.5-Flash-Thinking - Google DeepMind 2025.06.21
+Gemini-2.5-Pro - Google DeepMind 2025.03.25
+GPT-5 - OpenAI 2025.08.07
+Appendix D Experimental result on VisuLogic
+Table 7: Experimental result on VisuLogic Benchmark. The best results are marked bold and the second results are underlined.
+Models Param. Quantity Spatiality Position Attribute Style Other Overall
+Human - 45.3 52.7 71.1 50.0 47.5 44.2 51.4
 […]
-Human - 61.3 60.9 67.5 67.9 58.8 - - 61.9 63.0
-Open-Source MLLMs
-Minicpm-V-2.6 8B 22.2 23.1 38.9 23.1 27.4 1.0 11.0 28.2 21.9
-InternVL2.5-8B 8B 23.0 29.9 27.8 31.7 28.0 0.0 8.0 31.0 22.7
-InternVL2.5-8B-MPO 8B 25.7 29.9 20.0 27.9 27.4 0.0 10.0 31.0 22.4
-Deepseekvl2 27B 28.4 29.1 22.2 22.1 24.2 0.0 8.0 18.3 21.0
-InternVL2.5-38B 38B 22.6 26.5 28.9 27.9 26.8 1.0 9.0 31.0 21.9
-InternVL2.5-38B-MPO 38B 24.5 23.1 33.3 29.8 27.4 0.0 8.0 25.4 22.1
-Qwen2.5VL-32B 32B 25.7 23.9 24.4 29.8 28.7 1.0 9.0 25.4 22.1
-InternVL2.5-78B 78B 21.8 31.6 30.0 27.9 26.8 0.0 15.0 22.5 22.3
-InternVL2.5-78B-MPO 78B 26.8 31.6 25.6 26.9 24.8 0.0 13.0 23.9 22.7
-InternVL2.5-78B-MPO(cot) 78B 26.1 24.8 27.8 33.7 21.7 0.0 16.0 21.1 22.2
-Qwen2.5VL-72B 72B 24.5 23.1 23.3 29.8 22.3 2.0 12.0 29.6 21.3
-Qwen2.5VL-72B(cot) 72B 26.4 25.6 24.4 32.7 26.8 0.0 18.0 31.0 23.7
-Commercial products
-GPT-4o - 24.5 22.2 28.9 25.0 22.3 0.0 17.0 25.4 21.2
-GPT-4o(cot) - 28.7 28.2 34.4 27.9 22.9 0.0 11.0 33.8 23.9
-GPT-o1 - 31.0 28.2 33.3 25.0 26.1 0.0 15.0 26.8 24.5
-claude-3-7-sonnet - 25.3 23.1 30.0 26.0 25.5 1.0 12.0 28.2 22.0
-claude-3-7-sonnet-thinking - 21.1 26.5 41.1 30.8 26.8 9.0 13.0 26.8 23.8
-Gemini-2.0-flash-thinking - 17.6 12.8 35.6 26.9 17.2 10.0 11.0 25.4 18.7
-Gemini2.5-pro - 26.4 32.5 44.4 29.8 23.6 15.0 14.0 33.8 26.8
+Deepseek-R1 670B 27.7 23.5 24.0 27.8 23.0 35.0 26.6
+Qwen2.5-72B-Instruct 72B 30.2 24.4 27.5 26.5 26.8 30.8 28.0
+Claude-3.7-Sonnet (20250219) - 26.6 22.5 25.0 28.0 25.6 30.6 25.9
+Doubao-1.5-Pro-32k (20250115) - 30.0 22.5 25.0 25.6 30.0 24.1 26.6
+Close Source MLLMs
+GPT-4o-mini (20240718) - 27.2 23.4 23.5 18.3 31.1 16.7 24.3
+GPT-4o (20240806) - 28.6 24.7 27.2 26.8 20.0 25.9 26.3
+Kimi-latest - 24.9 29.4 26.5 28.0 16.7 26.9 25.9
+Doubao-1.5-Vision-Pro-32k (20250115) - 28.1 23.8 29.1 25.1 32.1 35.0 28.1
+Gemini-2.0-Pro (20250205) - 29.7 24.2 27.9 30.5 22.2 33.3 28.0
+Claude-3.7-Sonnet (20250219) - 22.7 27.3 27.9 28.0 22.2 22.2 24.8
+Open Source MLLMs
+LLaVA-v1.5 7B 26.1 24.2 23.5 17.1 31.1 22.2 24.6
+LLaVA-OneVision (SI) 7B 22.4 27.3 33.1 23.2 25.6 22.2 25.3
+Ovis2 8B 26.1 23.8 27.2 28.0 25.6 24.1 25.6
+Qwen2.5-VL-7B-Instruct 7B 27.6 20.9 25.2 23.2 37.8 25.0 26.0
+Qwen2.5VL-72B-Instruct 72B 25.2 23.8 27.2 25.6 25.6 34.3 26.2
+InternVL2.5-38B 8B 24.4 26.4 27.2 23.2 25.6 26.9 25.5
+InternVL2.5-78B 78B 26.6 26.0 26.5 26.8 31.1 30.6 27.3
+InternVL3-38B 38B 28.7 27.6 26.1 21.4 23.9 28.5 27.1
+InternVL3-78B 78B 27.7 26.1 31.6 26.3 21.3 32.3 27.7
+Qwen2.5-VL-7B-Instruct-SFT (VisuLogic) 7B 24.4 26.4 27.2 23.2 25.6 26.9 25.5
+Qwen2.5-VL-7B-Instruct-RL (VisuLogic) 7B 26.6 33.8 29.4 23.2 18.9 29.6 28.0
+InternVL2.5-38B-SFT (VisuLogic) 38B 30.6 29.4 20.6 25.6 30.0 25.0 27.9
+InternVL2.5-38B-RL (VisuLogic) 38B 31.2 31.2 26.5 30.5 30.0 38.9 31.1
 Ours
-Baseline(Qwen2.5VL-7B) 7B 23.4 28.2 28.9 23.1 21.7 0.0 14.0 23.9 20.9
-PRS-Qwen2.5VL-7B 7B 27.2 35.7 41.4 38.3 30.3 60.0 95.0 21.3 43.7
-4.2 Main Results
-Table 3 presents performance of various MLLMs on VisuRiddles benchmark. We summarize the key observations as follows:
-Performance of Open-Source MLLMs.
-Most open-source MLLMs demonstrate limited performance on VisuRiddles, with accuracies across core categories often close to random choice and consistently failing on the two high-level reasoning tasks, which reveals a substantial gap between current model capabilities and human performance in AVR.
-Performance of Commercial Models with Thinking Mode.
-Several commercial MLLMs with ‘thinking’ mode, such as GPT-o1, Claude-3-7-Sonnet-Thinking, and Gemini2.5-Pro, achieve comparatively improved performance on VisuRiddles.
-Their advantage is particularly notable in structured tasks like Sudoku and Raven.
-These results indicate that AVR can benefit from advanced reasoning capabilities.
-However, the overall performance remains unsatisfactory, suggesting that reasoning enhancements alone are insufficient to address the core challenges of this task.
-Effectiveness of Model Scaling and CoT Prompting.
-Strategies such as increasing model size and applying CoT prompting, though commonly used to enhance reasoning, are ineffective for AVR.
-Models with larger parameter scales do not consistently outperform smaller ones, and CoT yields only limited improvements on select tasks.
-These findings indicate that scaling in model parameter or inference time is insufficient to address the core challenges of AVR.
-Effectiveness of Our Perception-Enhanced Model.
-Our perception-enhanced model significantly outperforms other MLLMs across all categories of VisuRiddles, particularly in high-level reasoning tasks, Sudoku reasoning and Raven reasoning, demonstrating the model’s ability to handle structurally complex visual patterns.
-Importantly, these gains are achieved without changing model architecture or size, but solely through training on data with perceptual description, highlighting that the improvement arises from enhanced perceptual abstraction rather than stronger reasoning capacity.
-These results affirm the critical role of fine-grained perceptual supervision in abstract visual reasoning and emphasize the need to model intermediate perceptual processes for better performance on structurally complex tasks.
-Figure 4: Case study comparing different reasoning strategies on a VisuRiddles example.
-(a) QVQ-72B reflects a flawed loop under the ”thinking” mode.
-(b) Qwen2.5-VL-72B with CoT prompting exhibits incorrect perceptual understanding.
-(c) Our perception-enhanced model generates structured reasoning grounded in accurate perceptual abstraction, yielding the correct answer.
-4.3 Visualization Results
-To further analyze the reason that MLLMs underperform on AVR task, we examine two widely adopted reasoning enhancement strategies: inference-time scaling and CoT prompting.
-We conduct experiments on QVQ-72B-Preview (equipped with a ”think” mode), Qwen2.5-VL-72B (with CoT prompting), and our perception-enhanced model Qwen2.5-8B.
-Fig.
-4 visualizes a representative case of a broadly observed phenomenon uncovered during our experiments.
-The results reveal clear differences in the reasoning behaviors across these models.
-The inference-time scaling approach, which leverages ”think”-style prompting without any perceptual grounding, frequently leads to verbose yet logically inconsistent reasoning, often ending in incorrect answers.
-The CoT prompting, while occasionally yielding the correct answer, fails to guide the model along a valid reasoning path due to inaccurate visual perception.
-In such cases, correct predictions can be largely attributed to the single-choice format rather than abstract visual understanding.
-This observation aligns with CoT-based models perform close to random guessing on basic inference tasks, but struggle significantly on high-level inference tasks, shown in Tab.
-3.
-In contrast, our approach leverages fine-grained perception to produce concise, visually grounded reasoning and accurate answers.
-These results underscore the importance of perception in abstract visual reasoning.
-Inference-time scaling and CoT prompting alone cannot compensate for the lack of perceptual grounding, which often leads to incorrect or inconsistent reasoning.
-In contrast, introducing fine-grained perceptual supervision enables the model to build accurate and visually aligned reasoning paths, supporting reliable performance on complex visual tasks.
-4.4 Ablation Study
-Ablation Study on Bottlenecks in Abstract Visual Reasoning.
-To investigate whether the performance gap in abstract visual reasoning stems from perceptual limitations rather than reasoning ability, we conduct an ablation study using our perception-annotated synthetic dataset.
-We evaluate two representative MLLMs, GPT-4o (closed-source) and Qwen2.5VL-72B (open-source), on a subset of our synthetic dataset under two input settings: (i) raw abstract graphics (V), and (ii) structured perceptual descriptions (P).
-The experimental results in Tab.
-4 show that both models perform poorly when given only raw visual inputs, across both the five core perceptual dimensions and the two high-level reasoning tasks.
-Taking Sudoku reasoning as an example, this task requires recognizing digit values and their positions within a grid, and inferring a globally consistent solution.
-Under the Visual setting, both GPT-4o and Qwen2.5VL-72B fail to handle such constraint-based reasoning, while in the perception setting, accuracy increases from near-zero to 15% and 65%, respectively.
-This improvement primarily stems from structured perceptual descriptions compensating for the MLLMs’ limited perceptual abstraction capabilities.
-The difficulty stems from the nature of Sudoku images: small digits, dense layout, and absent semantic cues make them hard for MLLMs to interpret, though humans process such structure effortlessly through rapid visual scanning.
-Table 4: Performance Impact of Perceptual Descriptions vs.
-Visual Inputs in MLLMs.
-Where ”V” denotes inputting abstract graphics directly, while “P” refers to replacing graphics with structured perceptual descriptions.
-Model Num Styl Attr Posit Spat Sudo Rav Avg
-GPT-4o(V) 35.0 32.0 38.0 36.0 32.0 0.0 20.0 27.6
-GPT-4o(P) 62.0 (+27.0) 53.0 (+21.0) 80.0 (+42.0) 68.0 (+32.0) 100.0 (+68.0) 15.0 (+15.0) 25.0 (+5.0) 60.1(+32.5)
-Qwen2.5VL(V) 41.0 43.0 50.0 32.0 40.0 0.0 10.0 30.9
-Qwen2.5VL(P) 73.0(+32.0) 83.0(+40.0) 80.0(+30.0) 79.0(+47.0) 100.0(60.0) 65.0(+65.0) 35.0(+25.0) 73.6(+42.7)
-Table 5: Ablation study on dataset setting
-Model Num Styl Attr Posit Spat Sudo Rav Other Avg
-Baseline 23.4 28.2 28.9 23.1 21.7 0.0 14.0 23.9 20.9
-+20K 24.5 (+1.1) 29.8 (+1.6) 31.2 (+2.3) 26.1 (+3.0) 23.8 (+2.1) 19.0 (+19.0) 40.0 (+26.0) 22.0 (-1.9) 26.4 (+5.5)
-+40K 25.9 (+2.5) 31.8 (+3.6) 34.7 (+5.8) 29.7 (+6.6) 26.4 (+4.7) 30.0 (+30.0) 66.0 (+52.0) 23.5 (-0.4) 30.9 (+10.0)
-+60K 26.7 (+3.3) 33.8 (+5.6) 37.8 (+8.9) 33.5 (+10.4) 28.6 (+6.9) 53.0 (+53.0) 87.0 (+73.0) 24.8 (+0.9) 37.2 (+16.3)
-Full data 27.2 (+3.8) 35.7 (+7.5) 41.4 (+12.5) 38.3 (+15.2) 30.3 (+8.6) 60.0 (+60.0) 95.0 (+81.0) 21.3 (-2.6) 43.7 (+22.8)
-Ablation Study on Perceptually Annotated Synthetic Data.
-To evaluate the effectiveness of our synthesized data in enhancing model performance on abstract visual reasoning, we conduct a series of ablation studies with incremental training data.
-The results are shown in Tab.
-5.
-Increasing the amount of synthetic data with perceptual annotations leads to consistent improvements across all dimensions except for the ”Other” category, demonstrating that perceptually structured supervision provides effective guidance for internalizing abstract visual patterns that are difficult to learn from isolated answer-based supervision alone.
-This further highlights the importance of modeling intermediate perceptual processes in abstract visual reasoning.
-5 Conclusion
-In this work, we introduce VisuRiddles, a benchmark that systematically evaluates MLLMs’ AVR abilities across diverse, structurally rich tasks, along with an automated data synthesis framework that generates controllable, perceptually annotated reasoning problems.
-Experimental results demonstrate that even state-of-the-art MLLMs perform only marginally above random chance on VisuRiddles, exhibiting a substantial gap compared to human performance.
-Our analysis reveals that their limitations stem from inadequate perceptual abstraction rather than reasoning capacity, and that training with our synthetic data improves performance and interpretability.
-In sum, VisuRiddles enables fine-grained evaluation of reasoning, and our framework offers scalable data support for future training.
-More importantly, these findings underscore a central insight: enhancing perceptual abstraction is critical to advancing abstract visual reasoning in MLLMs.
+PAVR 7B 28.9 30.3 31.6 34.1 30.0 33.3 30.6